Please help
A tank holds 30 liters of a chemical solution. Currently, the solution has a strength of 60%. How much of this should be drained and replaced with a 90% solution to get a final strength of 80%?
How many liters should be drained and replaced with 90 % solution

Answer by stanbon(75887)

A tank holds 30 liters of a chemical solution. Currently, the solution has a strength of 60%. How much of this should be drained and replaced with a 90% solution to get a final strength of 80%?
---
Equation:
active - active + active = active
0.60*30 - 0.60x + 0.90x = 0.80*30
---------------------------------------
Multiply thru by 100 to get:
60*30 - 60x + 90x = 80*30
-----------------------------
30x = 20*30
x = 20 liters (amt of 60% to be replaced by 90% solution)
========================
Cheers,
stan H.